OpenStudy (anonymous):

determine where the graph is concave up or down y=-x^3+3x^2-2

OpenStudy (campbell_st):

well lots of ways to look at this easiest if to find the derivative... are you able to do that..?

OpenStudy (anonymous):

derivative is -3x^2+6x

OpenStudy (anonymous):

setting thhat equal to 0 I get x=0 and x=2

OpenStudy (campbell_st):

ok so they are the 2 stationary points... find the 2nd derivative and test the stationary points \[\frac{d^2y}{dx^2} > 0 ....minimum..... \frac{d^2y}{dx^2} < 0 .... maximum \]

OpenStudy (campbell_st):

\[\frac{d^2y}{dx^2} = -6x + 6\] so a point of inflection occurs when the 2nd derivative = 0 so a point of inflexion is x = 1

OpenStudy (campbell_st):

so what did you find...? by substituting x = 0 and x = 2 into the 2nd derivative...

OpenStudy (campbell_st):

so the curve has 2 parts... one is concave up... the other is concave down... than the change in concavity occurs at x = 1

OpenStudy (anonymous):

plugging in 0 I got 6, plugging in 2 I got -6

OpenStudy (campbell_st):

ok... so the curve is concave up at x = 0, concave down at x = 2  the change in concavity occurs at x = 1 hope this makes sense.